OpenStudy (anonymous):

A bumper car with a mass of 200 kilograms traveling at a velocity of 3.5 meters/second collides with a stationary bumper car with a mass of 250 kilograms. If the two bumper cars stick and move together at a velocity of 2.3 meters/second, what is the kinetic energy of the system after the collision?

OpenStudy (anonymous):

\[\color{green}{E_k = \frac{1}{2}(m_1 + m_2)v_f^2 \\ = \frac{1}{2}(450~\text{kg})(2.3~\text{m/s})^2}\] It appears that the additional information provided about the initial velocities is irrelevant.